# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ Abstract
Hyperspectral Image Classification (HSC) is a challenging task due to the high dimensionality and complex nature of Hyperspectral (HS) data. Traditional Machine Learning approaches while effective, face challenges in real-world data due to varying optimal feature sets, subjectivity in human-driven design, biases, and limitations. Traditional approaches encounter the curse of dimensionality, struggle with feature selection and extraction, lack spatial information consideration, exhibit limited robustness to noise, face scalability issues, and may not adapt well to complex data distributions. In recent years, DL techniques have emerged as powerful tools for addressing these challenges. This survey provides a comprehensive overview of the current trends and future prospects in HSC, focusing on the advancements from DL models to the emerging use of Transformers. We review the key concepts, methodologies, and state-of-the-art approaches in DL for HSC. We explore the potential of Transformer-based models in HSC, outlining their benefits and challenges. We also delve into emerging trends in HSC, as well as thorough discussions on Explainable AI and Interoperability concepts along with Diffusion Models (image denoising, feature extraction, and image fusion). Additionally, we address several open challenges and research questions pertinent to HSC. Comprehensive experimental results have been undertaken using three HS datasets to verify the efficacy of various conventional DL models and Transformers. Finally, we outline future research directions and potential applications that can further enhance the accuracy and efficiency of HSC. The Source code is available at https://github.com/mahmad00/Conventional-to-Transformer-for-Hyperspectral-Image-Classification-Survey-2024.
@@
Spatial-Spectral Feature; Hyperspectral Image Classification (HSC); Convolutional Neural Networks (CNNs); Spatial-Spectral Transformers (SSTs); Explainable AI and Interoperability; Diffusion Models.
@@ 1 Introduction
Hyperspectral (HS) Sensors capture detailed spectral information across a broad range of electromagnetic wavelengths [1]. Unlike traditional methods, Hyperspectral Images (HSIs) provide a continuous spectrum through numerous narrow bands, enabling precise material characterization and valuable Earth surface information extraction [2, 3]. Its significance in remote sensing lies in overcoming the limitations of multispectral imaging, allowing discrimination of subtle differences in material properties, and facilitating accurate land cover classification [4, 5].
Figure 1: Various Real-World Applications of HSI: This figure illustrates the diverse applications of HSI across different fields. In agriculture, HSI helps monitor crop health, detect diseases, and assess soil properties. In environmental monitoring, it aids in tracking pollution, mapping ecosystems, and assessing water quality. In the field of mineralogy, HSI is used for mineral identification and mining exploration. In healthcare, it assists in diagnosing diseases and monitoring tissue health. Additionally, HSI is employed in surveillance and security for target detection and material identification. Each application leverages the unique spectral signatures captured by HSI to provide detailed insights and improve decision-making processes.
@@
HSI’s key advantage lies in its distinguishing materials with similar visual appearances (e.g., color, texture, shape, structure, etc.) but distinct spectral properties [6]. This specificity is crucial in various applications, such as agriculture, forestry, urban planning, environmental monitoring, mineral exploration, disaster management, bloodstain identification, meat processing, contamination detection in red chili, and bakery products [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Figure 1 illustrates some HSI applications.
@@
In agriculture [7], HSI is used to monitor crop health, detect diseases, and estimate vegetation properties like chlorophyll content, optimizing irrigation and pest control for enhanced crop management. In forestry, HSI is used to assess forest health and species composition, and detects pests and diseases early, aiding conservation efforts. In geological studies [9, 10, 13] and mineral exploration [11, 12], HSI used to identify mineral deposits through unique spectral signatures, supporting resource exploration, mining, and environmental impact assessments. Additionally, HSI is used to assist urban planning by identifying land cover types, analyzing urban heat islands, monitoring urban expansion, and assessing environmental quality for sustainable development and infrastructure planning [24, 25].
@@
In short, HSI provides detailed spectral information, enhancing understanding of the Earth’s surface and supporting land cover classification, environmental assessment, change monitoring, and decision-making [26]. As technology advances and HS sensors become more accessible, HSI will play a vital role in addressing environmental challenges and facilitating sustainable resource management. This survey focuses on the field of Hyperspectral Image Classification (HSC), which has seen significant advancements, especially with the rise of Deep Learning (DL) models [27, 28]. The survey aims to comprehensively overview commonly used DL techniques for HSC. It will address primary challenges in HSC, not effectively tackled by Traditional Machine Learning (TML) methods, and emphasize how DL improves HSC performance.
1.1 Limitations of Existing Surveys
@@
However, existing surveys [29, 30, 1, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] on HSC have some shortcomings, including Limited Coverage: Some surveys may focus only on specific aspects of HSC, such as traditional methods or DL approaches, potentially overlooking other important aspects or emerging trends in the field. Outdated Information: Surveys published several years ago may not capture the latest advancements and breakthroughs in HSC techniques, rendering them less relevant for current research and applications. Lack of Depth: Some surveys may provide superficial coverage of the subject matter, lacking in-depth analysis or critical evaluation of different methodologies, techniques, and challenges in HSC. Narrow Perspective: Certain surveys may adopt a narrow perspective, focusing primarily on a single application domain or a particular subset of HSC techniques, thereby limiting their applicability to broader contexts. Addressing these shortcomings could enhance the effectiveness and utility of surveys on HSC, ensuring they provide comprehensive, up-to-date, and accessible resources for researchers and practitioners in the field.
@@
The current survey on HSC offers several advantages compared to existing survey papers: Comprehensive Coverage: The survey provides a comprehensive overview of current trends and future prospects in HSC, covering both traditional DL models and the emerging use of transformer-based models. This broad coverage ensures that readers gain insights into the full spectrum of approaches in the field. Focus on Emerging Technologies: By delving into the potential of Transformer-based models, the survey stays up-to-date with the latest advancements in DL for HSC. This focus on emerging technologies ensures that readers are informed about cutting-edge developments and potential future directions in the field. Thorough Discussions on Key Concepts: The survey includes thorough discussions on key concepts, methodologies, and state-of-the-art approaches in DL for HSC (for instance, RML, Convolutional Neural Network (CNN), and its variants, Deep Belief Network (DBN), Recurrent Neural Network (RNN), Autoencoders (AEs), Transformers, Explainable AI and Interoperability, and Diffusion Models). This depth of analysis helps readers understand the underlying principles and techniques used in the field. Inclusion of Experimental Results: The survey presents comprehensive experimental results conducted using three Hyperspectral (HS) datasets, verifying the efficacy of various DL models and transformers. This empirical evidence adds credibility to the survey’s findings and helps readers assess the practical implications of different approaches. Open Challenges and Future Directions: The survey addresses several open challenges and research questions pertinent to HSC, providing valuable insights into areas where further research is needed. Additionally, outlining future research directions and potential applications helps guide researchers toward promising avenues for exploration. Accessibility of Source Code: The availability of the survey’s source code on GitHub enhances transparency and reproducibility, allowing readers to access and validate the experimental results presented in the paper.
@@
1.2 Significance of Accurate Classification
Accurate classification in HSI analysis is crucial due to the wealth of complex information within the data’s numerous spectral bands. HSIs provide unique details about observed objects, making precise classification essential for diverse applications. The significance of accurate classification lies in extracting meaningful information such as land cover types and land-use patterns [41]. HSC techniques identify and map different land cover classes, aiding land management, resource planning, and environmental assessments [42, 43]. This accuracy is crucial for monitoring changes in land cover, providing valuable insights into deforestation, urban expansion, and ecosystem degradation, thereby supporting conservation efforts and sustainable land-use practices [44].
Accurate classification is crucial for environmental monitoring and assessment [10]. HS imagery detects factors like water quality, pollution sources, and ecosystem health, allowing identification of areas of concern, tracking changes, and evaluating environmental management strategies. This information is vital for ecosystem health, biodiversity conservation, and mitigating environmental risks. In agriculture, precise HSC supports crop monitoring, disease detection, and yield estimation [7]. It identifies stressed vegetation, enabling targeted interventions like irrigation or pesticide application. HSC also aids precision agriculture by providing information on soil composition, nutrient levels, and plant health, optimizing resource allocation for improved productivity.
In mineral exploration, precise classification of HS data identifies and maps mineral deposits and geological formations [11, 12]. Distinct spectral signatures of minerals aid prospecting and resource estimation, crucial for sustainable resource management, reducing exploration costs, and minimizing environmental impacts. Urban planning and infrastructure development benefit from accurate HSC [9]. Accurate classification of land cover types and urban features enables assessing urban growth patterns, monitoring changes, and planning for future development. It supports urban heat island analysis, transportation planning, and identifying green spaces, contributing to sustainable and livable cities.
Table 1: Structure of the Survey Paper
Section Topics
1 Brief Introduction, Limitation of Existing Surveys, Significance of Accurate Classification, and Brief Introduction to Deep Learning (DL)
2 Hyperspectral (HS) Data Representation
3 Learning Categories
4 Overview of Conventional Methods and limitations
5 Advantages of DL
6 Convolutional Neural Network (CNN), Spectral CNN, Spatial CNN, Spatial-Spectral CNN, Graph CNN, Future Research Directions
7 Deep Belief Network (DBN) and Future Research Directions
8 Recurrent Neural Network (RNN) and Future Research Directions
9 Autoencoders (AEs) and Future Research Directions
10 Limitations of Traditional DL Methods
11 Detailed Discussion on Transformers and its variants, Benefits, and challenges of Transformers, Emerging Trends, and Advantages
12 Explainable AI and Interoperability
13 Diffusion Models, their pros and cons for HS Imaging
14 Open Challenges and Research Questions
15 Experimental Results and Discussion
16 Conclusions and Future Research Directions
@@
1.3 Brief Introduction to Deep Learning
@@
DL, for instance, CNNs, has revolutionized HSC, outperforming conventional methods by leveraging neural networks and attention mechanisms [45, 46, 47]. CNNs excel in learning hierarchical representations and capturing spatial and spectral dependencies within HS data [48]. Their ability to automatically extract discriminative features makes them well-suited for accurate classification in HSI [49, 50]. RNNs, designed for sequential data, show promise in HSC for temporal analysis [51]. Exploiting the temporal dimension, RNNs capture contextual information for tasks like tracking land cover changes or monitoring dynamic environmental processes. Autoencoders (AEs), a neural network type, find application in HSC [52]. Comprising an encoder and decoder network, AEs compress input data into a low-dimensional representation (latent space) and attempt to reconstruct the original input [53]. In HSI, AEs are used for unsupervised feature learning and dimensionality reduction [54]. Training AEs on unlabeled HS data enables effective feature extraction and classification by capturing the underlying data structure [55].
@@
More recently, Transformers have gained significant attention in various domains, including natural language processing and computer vision [12, 49, 56, 4]. Transformers are based on a self-attention mechanism that allows the model to capture dependencies between different positions within a sequence. In the context of HSC, Transformers have shown promise in learning long-range dependencies and capturing non-local spectral relationships. By modeling the interactions between different spectral bands, transformers can effectively exploit the rich spectral information in HS data. This has led to improved classification performance and the ability to capture complex spectral patterns. Transformers can handle variable-length input sequences, making them suitable for HSIs of different spatial dimensions. Additionally, transformers offer interpretability, as the attention mechanism allows for visualizing the importance of different spectral bands in the classification process. This interpretability can be valuable for understanding the reasoning behind the model’s predictions and enhancing domain knowledge. In short, DL models, including CNNs, RNNs, AEs, and Transformers, demonstrate substantial potential in HSC. Leveraging their capacity to learn intricate representations, capture spatial and spectral dependencies, and exploit temporal information, these models contribute to significant advancements in accurate HSC, enhancing our understanding of the Earth’s surface across diverse applications.
1.4 Structure of the Survey
@@
The following sections are organized as follows (expressed in Table 1): Section 2 provides an in-depth description of the HS dataset, covering spectral, spatial, and spectral-spatial representation. In Section 3, various learning mechanisms for HSI processing are outlined. Section 4 examines TML for HSC along with their limitations. Section 5 explores the advantages of employing DL for HSC. Section 6 elaborates on CNN models for spectral, spatial, and spectral-spatial models along with Graph CNN models. Section 7 outlines DBN for HSC and potential research directions. Section 8 discusses RNNs for HSC and potential research directions. Section 9 details AEs for HSC and potential research directions. Section 10 draws limitations of traditional DL approaches. In Section 11, an extensive discussion on state-of-the-art deep models, specifically Transformers, includes a comparison with traditional deep models, benefits, and challenges for HSI analysis. Subsequently, Section 11 explores emerging trends and advancements in state-of-the-art models for HSI analysis. Section 12 discusses Explainable AI and Interoperability in HSC. Section 13 briefly describes the Diffusion models employed for HS imaging, while Section 14 outlines challenges and research questions. In Section 15, detailed experimental results for traditional DL models and Transformer-based deep models are presented. Finally, Section 16 concludes the survey and suggests immediate research directions.

### v4

@@ Abstract
Hyperspectral Image Classification (HSC) presents significant challenges owing to the high dimensionality and intricate nature of Hyperspectral (HS) data. While traditional Machine Learning (TML) approaches have demonstrated effectiveness, they often encounter substantial obstacles in real-world applications, including the variability of optimal feature sets, subjectivity in human-driven design, inherent biases, and methodological limitations. Specifically, TML suffers from the curse of dimensionality, difficulties in feature selection and extraction, insufficient consideration of spatial information, limited robustness against noise, scalability issues, and inadequate adaptability to complex data distributions. In recent years, Deep Learning (DL) techniques have emerged as robust solutions to address these challenges. This survey offers a comprehensive overview of current trends and future prospects in HSC, emphasizing advancements from DL models to the increasing adoption of Transformer and Mamba Model architectures. We systematically review key concepts, methodologies, and state-of-the-art approaches in DL for HSC. Furthermore, we investigate the potential of Transformer-based models and the Mamba Model in HSC, detailing their advantages and challenges. Emerging trends in HSC are explored, including in-depth discussions on Explainable AI and Interoperability concepts, alongside Diffusion Models for image denoising, feature extraction, and image fusion. To substantiate the efficacy of various conventional DL models, Transformers, and the Mamba Model, comprehensive experimental results were conducted on three HS datasets, yielding notable accuracies: 99.94% on Pavia University, 99.41% on Indian Pines, and 99.97% on the University of Houston dataset. Additionally, we identify several open challenges and pertinent research questions in the field of HSC. Finally, we outline future research directions and potential applications aimed at enhancing the accuracy and efficiency of HSC. The Source code is available at https://github.com/mahmad00/Conventional-to-Transformer-for-Hyperspectral-Image-Classification-Survey-2024.
@@
Spatial-Spectral Feature; Hyperspectral Image Classification (HSC); Deep Learning Techniques; Convolutional Neural Networks (CNNs); Spatial-Spectral Transformers (SSTs); Explainable AI and Interoperability; Diffusion Models; State-Space Models (Mamba).
@@ 1 Introduction
Hyperspectral (HS) Sensors capture detailed spectral information across a broad range of electromagnetic wavelengths (Ahmad et al., 2021c, ). Unlike traditional methods, Hyperspectral Images (HSIs) provide a continuous spectrum through numerous narrow bands, enabling precise material characterization and valuable Earth surface information extraction (Hong et al., 2024b, ; Li et al., 2024d, ). Its significance in remote sensing lies in overcoming the limitations of multispectral imaging, allowing discrimination of subtle differences in material properties, and facilitating accurate land cover classification (Cao et al., 2024b, ; Neupane et al.,, 2021).
Figure 1: Various Real-World Applications of Hyperspectral Imaging (HSI): Agriculture: HSI is utilized to monitor crop health through stress detection and nutrient assessment. It can identify diseases at an early stage by analyzing spectral changes in plant reflectance. Additionally, HSI helps assess soil properties, including moisture content and organic matter levels, which are crucial for optimizing crop yields. Environmental Monitoring: In this domain, HSI plays a vital role in tracking pollution levels by analyzing spectral signatures of pollutants in air and water. It facilitates ecosystem mapping by providing detailed information about vegetation types and health. HSI also effectively assesses water quality by detecting harmful algal blooms and monitoring sediment levels. Mineralogy: HSI is employed in mineral identification by differentiating between mineral types based on spectral characteristics. This capability is invaluable in mining exploration, as it aids in locating valuable mineral deposits and optimizing extraction processes. Healthcare: In the medical field, HSI assists in disease diagnosis by identifying abnormal tissue characteristics. It is also utilized in monitoring tissue health, particularly in assessing the efficacy of treatments through non-invasive spectral analysis. Surveillance and Security: HSI enhances surveillance capabilities by enabling target detection through the analysis of specific material signatures. It is also used in material identification, allowing for better security measures by distinguishing between benign and hazardous materials.
@@
HSI’s key advantage lies in its distinguishing materials with similar visual appearances (e.g., color, texture, shape, structure, etc.) but distinct spectral properties (Benham et al.,, 2024). This specificity is crucial in various applications, such as agriculture (Shuai et al.,, 2024; Deng et al.,, 2024), forestry (Ma et al., 2024c, ), urban planning (Khorat et al.,, 2024; Janecki and Dzierzbicka-Głowacka,, 2024), environmental monitoring (Song et al., 2024b, ; Wu et al., 2019a, ), mineral exploration (Hong et al.,, 2018), disaster management, smart cities Aryal and Dutta, (2015), bloodstain identification (Butt et al.,, 2022; Zulfiqar et al.,, 2021), meat processing (Ayaz et al., 2020b, ; Ayaz et al., 2020a, ), contamination detection in red chili (Khan et al.,, 2021, 2020), and bakery products (Saleem et al.,, 2020), further details are presented in Figure 1.
@@
In agriculture (Shuai et al.,, 2024), HSI is used to monitor crop health, detect diseases, and estimate vegetation properties like chlorophyll content, optimizing irrigation and pest control for enhanced crop management. In forestry, HSI is used to assess forest health and species composition, and detects pests and diseases early, aiding conservation efforts. In geological studies (Khorat et al.,, 2024; Janecki and Dzierzbicka-Głowacka,, 2024; Song et al., 2024b, ) and mineral exploration (Peng et al.,, 2024; Ahmad et al., 2024d, ), HSI used to identify mineral deposits through unique spectral signatures, supporting resource exploration, mining, and environmental impact assessments. Additionally, HSI is used to assist urban planning by identifying land cover types, analyzing urban heat islands, monitoring urban expansion, and assessing environmental quality for sustainable development and infrastructure planning (Tian et al.,, 2021; Kim and Guldmann,, 2014).
@@
In short, HSI provides detailed spectral information, enhancing understanding of the Earth’s surface and supporting land cover classification, environmental assessment, change monitoring, and decision-making (Ahmad et al., 2020a, ). As technology advances and HS sensors become more accessible, HSI will play a vital role in addressing environmental challenges and facilitating sustainable resource management. This survey focuses on the field of Hyperspectral Image Classification (HSC), which has seen significant advancements, especially with the rise of Deep Learning (DL) models (Ahmad and Mazzara,, 2024; Ghous et al.,, 2024). The survey aims to comprehensively overview commonly used DL techniques for HSC. It will address primary challenges in HSC, not effectively tackled by Traditional Machine Learning (TML) methods, and emphasize how DL improves HSC performance.
1.a Limitations of Existing Surveys
@@
However, existing surveys (Guerri et al.,, 2024; Ahmad et al., 2021c, ) on HSC have some shortcomings, including Limited Coverage: Some surveys may focus only on specific aspects of HSC, such as traditional methods or DL approaches, potentially overlooking other important aspects or emerging trends in the field. Outdated Information: Surveys published several years ago may not capture the latest advancements and breakthroughs in HSC techniques, rendering them less relevant for current research and applications. Lack of Depth: Some surveys may provide superficial coverage of the subject matter, lacking in-depth analysis or critical evaluation of different methodologies, techniques, and challenges in HSC. Narrow Perspective: Certain surveys may adopt a narrow perspective, focusing primarily on a single application domain or a particular subset of HSC techniques, thereby limiting their applicability to broader contexts. Addressing these shortcomings could enhance the effectiveness and utility of surveys on HSC, ensuring they provide comprehensive, up-to-date, and accessible resources for researchers and practitioners in the field.
@@
The current survey on HSC offers several advantages compared to existing surveys: Comprehensive Coverage: The survey provides a comprehensive overview of current trends and future prospects in HSC, covering both traditional DL models and the emerging use of transformer-based models. This broad coverage ensures that readers gain insights into the full spectrum of approaches in the field. Focus on Emerging Technologies: By delving into the potential of Transformer-based models, the survey stays up-to-date with the latest advancements in DL for HSC. This focus on emerging technologies ensures that readers are informed about cutting-edge developments and potential future directions in the field. Thorough Discussions on Key Concepts: The survey includes thorough discussions on key concepts, methodologies, and state-of-the-art approaches in DL for HSC (for instance, RML, Convolutional Neural Network (CNN), and its variants, Deep Belief Network (DBN), Recurrent Neural Network (RNN), Autoencoders (AEs), Transformers, Explainable AI and Interoperability, and Diffusion Models). This depth of analysis helps readers understand the underlying principles and techniques used in the field. Inclusion of Experimental Results: The survey presents comprehensive experimental results conducted using three Hyperspectral (HS) datasets, verifying the efficacy of various DL models and transformers. This empirical evidence adds credibility to the survey’s findings and helps readers assess the practical implications of different approaches. Open Challenges and Future Directions: The survey addresses several open challenges and research questions pertinent to HSC, providing valuable insights into areas where further research is needed. Additionally, outlining future research directions and potential applications helps guide researchers toward promising avenues for exploration. Accessibility of Source Code: The availability of the survey’s source code on GitHub enhances transparency and reproducibility, allowing readers to access and validate the experimental results presented in the survey.
@@
1.b Significance of Accurate Classification
Accurate classification in HSI analysis is crucial due to the wealth of complex information within the data’s numerous spectral bands. HSIs provide unique details about observed objects, making precise classification essential for diverse applications. The significance of accurate classification lies in extracting meaningful information such as land cover types and land-use patterns (Xu et al.,, 2024). HSC techniques identify and map different land cover classes, aiding land management, resource planning, and environmental assessments (Ahmad,, 2021; Ahmad et al.,, 2017). This accuracy is crucial for monitoring changes in land cover, providing valuable insights into deforestation, urban expansion, and ecosystem degradation, thereby supporting conservation efforts and sustainable land-use practices (Patel et al.,, 2024).
Accurate classification is crucial for environmental monitoring and assessment (Janecki and Dzierzbicka-Głowacka,, 2024). HS imagery detects factors like water quality, pollution sources, and ecosystem health, allowing identification of areas of concern, tracking changes, and evaluating environmental management strategies. This information is vital for ecosystem health, biodiversity conservation, and mitigating environmental risks. In agriculture, precise HSC supports crop monitoring, disease detection, and yield estimation (Shuai et al.,, 2024). It identifies stressed vegetation, enabling targeted interventions like irrigation or pesticide application. HSC also aids precision agriculture by providing information on soil composition, nutrient levels, and plant health, optimizing resource allocation for improved productivity.
In mineral exploration, precise classification of HS data identifies and maps mineral deposits and geological formations (Peng et al.,, 2024; Ahmad et al., 2024d, ). Distinct spectral signatures of minerals aid prospecting and resource estimation, crucial for sustainable resource management, reducing exploration costs, and minimizing environmental impacts. Urban planning and infrastructure development benefit from accurate HSC (Khorat et al.,, 2024). Accurate classification of land cover types and urban features enables assessing urban growth patterns, monitoring changes, and planning for future development. It supports urban heat island analysis, transportation planning, and identifying green spaces, contributing to sustainable and livable cities.
@@
1.c Brief Introduction to Deep Learning
@@
DL, for instance, CNNs, has revolutionized HSC, outperforming conventional methods by leveraging neural networks and attention mechanisms (Ahmad et al., 2022c, ; Ji et al.,, 2024; Ghorbanzadeh et al.,, 2019). CNNs excel in learning hierarchical representations and capturing spatial and spectral dependencies within HS data (Ahmad et al., 2021b, ). Their ability to automatically extract discriminative features makes them well-suited for accurate classification in HSI (Ran et al.,, 2023; Ahmad et al., 2022a, ). RNNs, designed for sequential data, show promise in HSC for temporal analysis (Gong et al.,, 2024). Exploiting the temporal dimension, RNNs capture contextual information for tasks like tracking land cover changes or monitoring dynamic environmental processes. Autoencoders (AEs), a neural network type, find application in HSC (Gogineni et al.,, 2024). Comprising an encoder and decoder network, AEs compress input data into a low-dimensional representation (latent space) and attempt to reconstruct the original input (Ahmad et al., 2020c, ). In HSI, AEs are used for unsupervised feature learning and dimensionality reduction (Ahmad et al., 2019a, ). Training AEs on unlabeled HS data enables effective feature extraction and classification by capturing the underlying data structure (Ahmad et al., 2019c, ).
@@
More recently, Transformers have gained significant attention in various domains, including natural language processing and computer vision (Ahmad et al., 2024d, ; Ran et al.,, 2023; Jiao et al.,, 2023; Cao et al., 2024b, ). Transformers are based on a self-attention mechanism that allows the model to capture dependencies between different positions within a sequence. In the context of HSC, Transformers have shown promise in learning long-range dependencies and capturing non-local spectral relationships. By modeling the interactions between different spectral bands, transformers can effectively exploit the rich spectral information in HS data. This has led to improved classification performance and the ability to capture complex spectral patterns. Transformers can handle variable-length input sequences, making them suitable for HSIs of different spatial dimensions. Additionally, transformers offer interpretability, as the attention mechanism allows for visualizing the importance of different spectral bands in the classification process. This interpretability can be valuable for understanding the reasoning behind the model’s predictions and enhancing domain knowledge. In short, DL models, including CNNs, RNNs, AEs, and Transformers, demonstrate substantial potential in HSC. Leveraging their capacity to learn intricate representations, capture spatial and spectral dependencies, and exploit temporal information, these models contribute to significant advancements in accurate HSC, enhancing our understanding of the Earth’s surface across diverse applications.
1.d Structure of the Survey
@@
The following sections are organized as follows: Section 2 provides an in-depth description of the HS dataset, covering spectral, spatial, and spectral-spatial representation. In Section 3, various learning mechanisms for HSI processing are outlined. Section 4 examines TML for HSC along with their limitations. Section 5 explores the advantages of employing DL for HSC. Section 6 elaborates on CNN models for spectral, spatial, and spectral-spatial models along with Graph CNN models. Section 7 outlines DBN for HSC and potential research directions. Section 8 discusses RNNs for HSC and potential research directions. Section 9 details AEs for HSC and potential research directions. Section 10 draws limitations of traditional DL approaches. In Section 11, an extensive discussion on state-of-the-art deep models, specifically Transformers, includes a comparison with traditional deep models, benefits, and challenges for HSI analysis. Subsequently, Section 11 explores emerging trends and advancements in state-of-the-art models for HSI analysis. Section 12 discusses Explainable AI and Interoperability in HSC. Section 13 briefly describes the Diffusion models employed for HS imaging. Section 14 describes State Space Modeling (SSM)–Mamba for HSC, while Section 15 outlines challenges and research questions. In Section 16, detailed experimental results for traditional DL models and Transformer-based deep models are presented. Finally, Section 17 concludes the survey and suggests immediate research directions.